OpenStudy (keana):

Identify the domain of the graph of y = -x^2 - 6x - 13 A: All real numbers B: x ≤ -4 C: x ≥ -6 D: x ≥ -2

OpenStudy (anonymous):

this is a parabola that opens up vertex is at \(x=-\frac{b}{2a}=-\frac{-12}{2}=6\) second coordinate of the vertex is \[6^2-12\times 6+3=36-72+3=-36+3=-33\] range is \[y\geq -33\]

OpenStudy (anonymous):

so none of the above, unless there is a typo on your question

OpenStudy (anonymous):

damn satellite! you're a 100?!

OpenStudy (keana):

well that isn't possible . the equation IS right !

OpenStudy (anonymous):

the question is asking for the domain... it's all real numbers...